# Simplex-to-Euclidean Bijections for Categorical Flow Matching

### v1

@@ 1 INTRODUCTION
We study the problem of learning and generating samples from probability distributions supported on the unit simplex, a setting that naturally arises when working with categorical compositional data (vectors of non-negative components that sum to one). Learning distributions on the simplex provides a natural framework for many real-world applications. For example, in computational biology, sequences can be encoded as categorical variables e.g. for conditional generation of DNA sequences (Avdeyev et al., 2023) or proteins (Campbell et al., 2024), and compositional data naturally arises in geology, chemistry, design and economics (Aitchison, 1982; Chereddy and Femiani, 2025; Diederen and Zamboni, 2025).
@@ 3.1 Simplex-to-Euclidean bijections
Aitchison (1981, 1982) noted that for compositional data (vectors in the open simplex), the logratio between the components provides a principled method for constructing smooth bijections from the open simplex to Euclidean space. Two classical examples, which will serve as building blocks of the transforms we later propose, are the additive logratio (ALR), for $1\leq k\leq D$
@@
The naive ALR and MLR mappings could in principle be used as such, and recently Chereddy and Femiani (2025) indeed used ALR for modeling discrete data (for CAD generation) with diffusion models, without accounting for the order invariance in any way. We will next introduce two concrete mappings; one is invariant to the order, resolving the fundamental limitation, whereas the other improves on MLR by aligning it to the center point on each space.
@@
where $\bm{H}\in\mathbb{R}^{D\times K}$ a Helmert matrix (Lancaster, 1965) and the rows of $\bm{H}$ form an orthonormal basis that spans the tangent space of the simplex $T_{\bm{x}}\Delta^{D}=\{\bm{x}\in\mathbb{R}^{K}:\sum_{k=1}^{K}x_{k}=0\}$. Characterization of the Helmert matrix and other details are provided in the Supplement. The volume change is
@@
$\bm{x}$ $=\mathrm{mlr}(\bm{y}),\ y_{k}=z_{k}+\log\!\left(\frac{1}{K-k}\right).$
@@ 3.3 Training, sampling and evaluation
We cannot directly evaluate the categorical probability $\Pr(C=k)$, since the model only defines a continuous density on the interior. We do not need it for learning (training uses the interpolated points) or sampling (obtained by the $\arg\max$-operation) and it is not provided by many alternatives either. For example, Cheng et al. (2024) only provides an overly loose lower bound (see Supplement for a demonstration). For evaluation we hence need to estimate $\Pr(C=k)$.
@@ 4 EXPERIMENTS
We evaluate our approach on five tasks. We compare against other continuous models DirichletFM (Stark et al., 2024), DDSM (Avdeyev et al., 2023), Bit-Diffusion (Chen et al., 2023) and $\alpha$-Flow (Cheng et al., 2025). As additional baselines, we report results with some methods working directly in the discrete space, DFM (Gat et al., 2024), D3PM (Austin et al., 2021), MDLM (Sahoo et al., 2024), SEDD (Lou et al., 2024) and MultiFlow (Campbell et al., 2024), to measure the gap between the two families.
We also report results for a baseline using Euclidean geometry directly on the simplex (Chen and Lipman, 2024; Stark et al., 2024), denoting it by LinearFM. As explained in Section 2.1, our method is used with minibatch OT (w/OT) and without OT, we report both results. We fix $\lambda=\tfrac{1}{2}$ and $\alpha=100$ in all cases.
@@
The previous experiments all considered problems of fixed dimensionality. To study the performance as a function of $K$, we consider a setup where $10^{6}$ samples are drawn from discrete distributions of $K=2^{1},\dots,2^{9}$ categories, with some probability vector $\bm{p}\in\Delta^{D}$. After training the model, we draw $10^{5}$ samples and compute the KL divergence between the empirical density of the samples and the true distribution. Fig. 4 shows our method generally outperforms SFM and LinearFM especially for medium dimensionalities, and is comparable to the discrete-state SEDD for dimensionality up to $K=2^{7}$. The Supplement has additional experiments under the same setup on the effect of $\alpha$ and $\lambda$, which were kept fixed in all of the main experiments.
@@ 5 DISCUSSION
Diederen and Zamboni (2025) used the ILR transformation in modeling compositional data, as one transformation within a composition of multiple ones. They did not, however, consider discrete observations at all. Potapczynski et al. (2020), in turn, considered an alternative bijection building on $\mathrm{softmax}_{++}$ that also allows recovering discrete data with $\arg\max$, but their solution is designed specifically for use a variational autoencoder as the generative model, with no obvious generalization for other models.
Chereddy and Femiani (2025) proposed a diffusion model that leverages special cases of our machinery. Their Gaussian-Softmax diffusion model uses the push-forward of a Gaussian with the additive logratio transform (Eq. 3), using a deterministic interpolation scheme similar to ours ($\lambda\bm{e}_{k}+(1-\lambda)\tfrac{1}{K}$ with $\lambda=0.99$) to handle discrete observations. However, their transformation leaves the mass very close to the borders and the solution is specifically geared towards use of the order-dependent alr transformation. Cheng et al. (2025) proposed an alternative bijection on the logit-simplex space, as a limiting case of the information geometry, resulting in a geometry that is intrinsically different from ours. Closer inspection of the differences would be highly interesting future work. Mahabadi et al. (2024) and Tae et al. (2025) also used the logit-simplex space, but treated each discrete token as a point in the probability simplex close to a vertex, rather than making a formal connection with the vertices and well-defined points in the interior.
@@ A.1 Proof of proposition 4
A.2 Proof of proposition 5
Proposition 5.
Proof.
Because $\lambda\geq\tfrac{1}{2}$ and Prop. 4 the mixture places almost surely all mass of component $k$ inside $\mathcal{R}_{k}$, hence $p_{k}=\int_{\mathcal{R}_{k}}q_{\lambda}(\bm{x})\,d\bm{x}$. For any measurable set $B$, $\big|\int_{B}(\tilde{q}(\bm{x})-q_{\lambda}(\bm{x}))\,d\bm{x}\big|\leq\int_{B}|\tilde{q}(\bm{x})-q_{\lambda}(\bm{x})|\,d\bm{x}$, giving a point-wise bound. Summing and using $\sum_{k}\int_{\mathcal{R}_{k}}|\tilde{q}(\bm{x})-q_{\lambda}(\bm{x})|\,d\bm{x}\leq\int|\tilde{q}(\bm{x})-q_{\lambda}(\bm{x})|\,d\bm{x}$ yields
@@ A.3 Proof of proposition 6
Recall D:=K1D:=K-1 and $\bm{x}\sim\operatorname{Dir}(\alpha,..,\alpha)$. The first and second moments of the log-random variable are:
@@ C.2 Effect of the parameters
Using the same setup as in Scalability, we train a velocity network across the values $\alpha\in\{1,10,100,\infty\}$, $\lambda\in\{0.5,0.75,0.99\}$, with and without scaling (see Eq. 11). We only consider the ILR bijection. The case $\alpha=1$ corresponds to the uniform distribution over each $\arg\max$ region ($x_{i}>x_{j}$ for all $i\neq j$), and $\alpha=\infty$ to a deterministic interpolation, namely $\lambda\bm{c}+(1-\lambda)\tfrac{1}{K}$.
Figure 7 shows that scaling improves the accuracy when there are only 2 categories, but scaling degrades the performance once the number of categories exceeds $2^{5}$. The uniform distribution ($\alpha=1$) is slightly more accurate for $\lambda\in\{0.5,0.75\}$ than $\lambda=0.99$. Otherwise the performance of our method is largely insensitive to $\lambda$ or $\lambda$, with all trained vector fields exhibiting similar accuracy.
Figure 7: The KL divergence between the true and estimated categorical probabilities for $\alpha\in\{1,10,100,\infty\}$, $\lambda\in\{0.5,0.75,0.99\}$, with and without scaling.

### v2

@@ 1 INTRODUCTION
We study the problem of learning and generating samples from probability distributions supported on the unit simplex, a setting that naturally arises when working with compositional data (vectors of non-negative components that sum 1o one). Learning distributions on the simplex provides a natural framework for many real-world applications. For example, in computational biology, sequences can be encoded as categorical variables e.g. for conditional generation of DNA sequences (Avdeyev et al., 2023) or proteins (Campbell et al., 2024), and compositional data naturally arises in geology, chemistry, design and economics (Aitchison, 1982; Chereddy and Femiani, 2025; Diederen and Zamboni, 2025).
@@ 3.1 Simplex-to-Euclidean bijections
Atchison and Shen (1980); Aitchison (1981, 1982) noted that for compositional data (vectors in the open simplex), the logratio between the components provides a principled method for constructing smooth bijections from the open simplex to Euclidean space. Two classical examples, which will serve as building blocks of the transforms we later propose, are the additive logratio (ALR), for $1\leq k\leq D$
@@
The naive ALR and MLR mappings could in principle be used as such (see the Supplement C.4), and recently Chereddy and Femiani (2025) indeed used ALR for modeling discrete data (for CAD generation) with diffusion models, without accounting for the order invariance in any way. We will next introduce two concrete mappings; one is invariant to the order, resolving the fundamental limitation, whereas the other improves on MLR by aligning it to the center point on each space.
@@
where $\bm{H}\in\mathbb{R}^{D\times K}$ a Helmert matrix (Lancaster, 1965) and the rows of $\bm{H}$ form an orthonormal basis that spans the tangent space of the simplex $T_{\bm{x}}\Delta^{D}=\{\bm{x}\in\mathbb{R}^{K}:\sum_{k=1}^{K}x_{k}=0\}$. Characterization of the Helmert matrix, proof of $\mathcal{O}(K)$ complexity and other details are provided in the Supplement A.6. The volume change is
@@
$\bm{x}$ $=\mathrm{mlr}^{-1}(\bm{y}),\ y_{k}=z_{k}+\log\!\left(\frac{1}{K-k}\right).$
@@ 3.2 Handling discrete data
$\mathrm{TV}(\bm{p},\hat{\bm{p}})=\tfrac{1}{2}\sum_{k=1}^{K}|\hat{p}_{k}-p_{k}|\;\leq\;\tfrac{1}{2}\,\|\tilde{q}-q_{\lambda}\|_{1}.$
In particular, $\|\tilde{q}-q_{\lambda}\|_{1}\to 0$ implies $\hat{\bm{p}}\to\bm{p}$ in total variation, and the $\arg\max$ discretization of $\tilde{q}$ recovers exactly the true categorical distribution.
@@ 3.3 Training, sampling and evaluation
We cannot directly evaluate the categorical probability $\Pr(C=k)$, since the model only defines a continuous density on the interior. We do not need it for learning (training uses the interpolated points) or sampling (obtained by the $\arg\max$-operation) and it is not provided by many alternatives either. For example, Cheng et al. (2024) only provides an overly loose lower bound (see Supplement C.1 for a demonstration). Nonetheless, for evaluation it is useful to estimate $\Pr(C=k)$
@@ 4 EXPERIMENTS
We evaluate our approach on five tasks. We compare against other continuous models DirichletFM (Stark et al., 2024), DDSM (Avdeyev et al., 2023), Bit-Diffusion (Chen et al., 2023), Gumbel-Softmax FM (Tang et al., 2025) and $\alpha$-Flow (Cheng et al., 2025). As additional baselines, we report results with some methods working directly in the discrete space, DFM (Gat et al., 2024), D3PM (Austin et al., 2021), MDLM (Sahoo et al., 2024), SEDD (Lou et al., 2024) and MultiFlow (Campbell et al., 2024), to measure the gap between the two families.
We also report results for a baseline using Euclidean geometry directly on the simplex (Chen and Lipman, 2024; Stark et al., 2024), denoting it by LinearFM. As explained in Section 2.1, our method is used with minibatch OT (w/OT) and without OT, we report both results. We fix $\lambda=\tfrac{1}{2}$ and $\alpha=100$ in all cases. Our code implementation is available at github.com/williwilliams3/simplexfm.
@@
The previous experiments all considered problems of fixed dimensionality. To study the performance as a function of $K$, we consider a setup where $10^{6}$ samples are drawn from discrete distributions of $K=2^{1},\dots,2^{9}$ categories, with some probability vector $\bm{p}\in\Delta^{D}$. After training the model, we draw $10^{5}$ samples and compute the KL divergence between the empirical density of the samples and the true distribution. Fig. 4 shows our method generally outperforms SFM and LinearFM especially for medium dimensionalities, and is comparable to the discrete-state SEDD for dimensionality up to $K=2^{7}$. The Supplement C.2 has additional experiments under the same setup on the effect of $\alpha$ and $\lambda$, which were kept fixed in all of the main experiments.
@@ 5 DISCUSSION
Diederen and Zamboni (2025) used the ILR transformation in modeling compositional data, as one transformation within a composition of multiple ones. They did not, however, consider discrete observations at all. Potapczynski et al. (2020), in turn, considered an alternative bijection building on $\mathrm{softmax}_{++}$ that also allows recovering discrete data with $\arg\max$, but their solution is designed specifically for a variational autoencoder as the generative model, with no obvious generalization for other models.
Chereddy and Femiani (2025) proposed a diffusion model that leverages special cases of our machinery. Inspired by logistic-normal models (Atchison and Shen, 1980), their Gaussian-Softmax diffusion model uses the push-forward of a Gaussian with the additive logratio transform (Eq. 3), using a deterministic interpolation scheme similar to ours ($\lambda\bm{e}_{k}+(1-\lambda)\tfrac{1}{K}$ with $\lambda=0.99$) to handle discrete observations. However, their transformation leaves the mass very close to the borders and the solution is specifically geared towards use of the order-dependent ALR transformation. Cheng et al. (2025) proposed an alternative bijection on the logit-simplex space, as a limiting case of the information geometry, resulting in a geometry that is intrinsically different from ours. Closer inspection of the differences would be highly interesting future work. Mahabadi et al. (2024) and Tae et al. (2025) also used the logit-simplex space, but treated each discrete token as a point in the probability simplex close to a vertex, rather than making a formal connection with the vertices and well-defined points in the interior.
@@ Acknowledgements
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Yes]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
Simplex-to-Euclidean Bijections for Categorical Flow Matching:
@@ A.1 Proof of proposition 1
Proposition 1.
Proof.
Because $\lambda\geq\tfrac{1}{2}$ and Prop. 2 the mixture places almost surely all mass of component $k$ inside $\mathcal{R}_{k}$, hence $p_{k}=\int_{\mathcal{R}_{k}}q_{\lambda}(\bm{x})\,d\bm{x}$. For any measurable set $B$, $\big|\int_{B}(\tilde{q}(\bm{x})-q_{\lambda}(\bm{x}))\,d\bm{x}\big|\leq\int_{B}|\tilde{q}(\bm{x})-q_{\lambda}(\bm{x})|\,d\bm{x}$, giving a point-wise bound. Summing and using $\sum_{k}\int_{\mathcal{R}_{k}}|\tilde{q}(\bm{x})-q_{\lambda}(\bm{x})|\,d\bm{x}\leq\int|\tilde{q}(\bm{x})-q_{\lambda}(\bm{x})|\,d\bm{x}$ yields
A.2 Proof of proposition 2
@@ A.3 Proof of proposition 3
Recall D:=K1D:=K-1 and $\bm{x}\sim\operatorname{Dir}(\alpha,..,\alpha)$. Let $\psi$ denote the digamma function and $\psi^{\prime}=\frac{d}{dx}\psi(x)$ the trigamma function. The first and second moments of the log-random variable are:
@@ A.6 Isometric logratio transform
Computational complexity
We derive the linear computational complexity of the ILR bijection. The (non-full) Helmert matrix $\bm{H}\in\mathbb{R}^{D\times K}$ has a simple recursive structure. For $i=1,\dots,D$, the $i$-th row contains $i$ entries equal to $1/\sqrt{i(i+1)}$, one entry equal to $i/\sqrt{i(i+1)}$ in column i+1i{+}1, and zeros elsewhere:
$\bm{H}=\begin{bmatrix}\frac{1}{\sqrt{1\cdot 2}}&-\frac{1}{\sqrt{1\cdot 2}}&0&0&\cdots&0\\[6.0pt] \frac{1}{\sqrt{2\cdot 3}}&\frac{1}{\sqrt{2\cdot 3}}&-\frac{2}{\sqrt{2\cdot 3}}&0&\cdots&0\\[6.0pt] \frac{1}{\sqrt{3\cdot 4}}&\frac{1}{\sqrt{3\cdot 4}}&\frac{1}{\sqrt{3\cdot 4}}&-\frac{3}{\sqrt{3\cdot 4}}&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ \frac{1}{\sqrt{D\cdot K}}&\frac{1}{\sqrt{D\cdot K}}&\cdots&\frac{1}{\sqrt{D\cdot K}}&\frac{1}{\sqrt{D\cdot K}}&-\frac{D}{\sqrt{D\cdot K}}\end{bmatrix}$.
Let $\bm{v}=\log\bm{x}$. Then $\bm{z}=\bm{H}\bm{v}$ has components
$z_{i}=\frac{1}{\sqrt{i(i+1)}}\!\left(\sum_{j=1}^{i}v_{j}-i\,v_{i+1}\right).$
Defining the cumulative sum $S_{i}=\sum_{j=1}^{i}v_{j}$ (computed in $\mathcal{O}(K)$), we obtain
$z_{i}=\frac{S_{i}-i\,v_{i+1}}{\sqrt{i(i+1)}},\qquad i=1,\dots,D,$
Thus the full matrix–vector product $\bm{z}=\bm{H}\bm{v}$ can be computed in $\mathcal{O}(K)$ time without explicitly forming $\bm{H}$.
@@ Appendix B EXPERIMENTAL DETAILS
Table 5: Run time measurements with varying data dimensionality $D$ and fixed batch size $B=512$. All results are reported in milliseconds as a mean with two standard deviations across 1000 iterations.
Method $D=2^{2}$ (ms) $D=2^{3}$ (ms) $D=2^{4}$ (ms) $D=2^{5}$ (ms) $D=2^{6}$ (ms) $D=2^{7}$ (ms)
SFM $45\pm 3$ $45\pm 4$ $48\pm 3$ $48\pm 3$ $50\pm 3$ $54\pm 4$
FM-$\mathring{\Delta}$(ALR) $76\pm 12$ $74\pm 10$ $76\pm 12$ $2\pm 12 88$$8\pm 5$ $85\pm 10$
FM-$\mathring{\Delta}$(ILR) $81\pm 12$ $84\pm 12$ $83\pm 12$ $5\pm 12 88$$8\pm 6$ $92\pm 8$
Table 6: Run time measurements with varying batch size $B$ and fixed data dimensionality $D=64$. All results are reported in milliseconds as a mean with two standard deviations across 1000 iterations.
Method $B=2^{6}$ (ms) $B=2^{7}$ (ms) $B=2^{8}$ (ms) $B=2^{9}$ (ms) $B=2^{10}$ (ms) $B=2^{11}$ (ms)
SFM $15\pm 2$ $18\pm 1$ $29\pm 1$ $51\pm 3$ $93\pm 10$ $183\pm 36$
FM-$\mathring{\Delta}$(ALR) $13\pm 4$ $18\pm 6$ $42\pm 6$ $89\pm 6$ $171\pm 18$ $346\pm 36$
FM-$\mathring{\Delta}$(ILR) $13\pm 4$ $18\pm 6$ $43\pm 9$ $83\pm 12$ $166\pm 12$ $339\pm 32$
@@ C.2 Effect of the parameters
Using the same setup as in Scalability, we train a velocity network across the values $\alpha\in\{1,10,100,\infty\}$, $\lambda\in\{0.5,0.75,0.99\}$, with and without scaling (see Eq. 11). We only consider the ILR bijection. The case $\alpha=1$ corresponds to the uniform distribution over each $\arg\max$ region ($x_{i}>x_{j}$ for all $i\neq j$), and $\alpha=\infty$ to a deterministic interpolation, namely $\lambda\bm{c}+(1-\lambda)\tfrac{1}{K}$. We generate $100,000$ samples $5$ times and plot the mean values with 2 standard deviation bands.
Figure 7 shows that scaling improves the accuracy when there are only 2 categories, but scaling degrades the performance once the number of categories exceeds $2^{5}$. The uniform distribution ($\alpha=1$) is slightly more accurate for $\lambda\in\{0.5,0.75\}$ than $\lambda=0.99$. The performance of the method is more sensitive to the the values of $\alpha$ and $\lambda$ as $K$ increases, but we intentionally did not attempt fine-tuning $\alpha$ and $\lambda$ to maximize the empirical performance, since we want an easy-to-use pipeline.
Figure 7: The KL divergence between the true and estimated categorical probabilities for $\alpha\in\{1,10,100,\infty\}$, $\lambda\in\{0.5,0.75,0.99\}$, with and without scaling. The bands are two standard deviations away from the mean.
C.3 Run Times
We compare the run times between SFM and FM-$\mathring{\Delta}$(ALR) with the same setup as the Scalability experiment. To measure the effect of the number of categories, we consider dimensions $D=2^{2},\ldots,2^{7}$ with a fixed batch size of $B=512$, presented in Table 5. Then, to measure the effect of batch size, we consider batch sizes $B=2^{6},2^{7},\ldots,2^{1}1$ with a fixed $D=64$, presented in Table 6. All the run times are reported with mean and two standard deviations over 1000 training iterations in machines with Xeon E5 2680 v3 2.50GHz CPUs, we do not use GPUs for these measurements.
C.4 Numerical results for Additive Logratio and Multiplicative Logratio bijections
In Section 3.1 we defined ALR and MLR as possible bijections, before introducing our proposed bijections SB and ILR. It would still be useful to clarify how well ALR and MLR would work.
We now ran the Checkerboard, BMNIST and DNA experiments with the ALR and MLR mappings, reporting the results in Table 7. SB remains the best on DNA and Checkerboard and ILR on binarized MNIST, but both ALR and MLR also work relatively well. SB and MLR are highly similar (for BMNIST with 2 classes they are exactly the same map), as they should because the former mostly just stabilizes the computation by centering.
Table 7: Numerical results for ALR and MLR bijections.
Method Checkerboard BMNIST DNA
FM-$\mathring{\Delta}$(ALR) 6.4% 4.98 0.029
FM-$\mathring{\Delta}$(ALR)w/OT 5.22 0.037
FM-$\mathring{\Delta}$(MLR) 5.7% 4.93 0.027
FM-$\mathring{\Delta}$(MLR)w/OT 4.51 0.024
FM-$\mathring{\Delta}$(SB) 5.4% 4.93 0.028
FM-$\mathring{\Delta}$(SB)w/OT 4.51 0.021
FM-$\mathring{\Delta}$(ILR) 6.8% 4.36 0.026
FM-$\mathring{\Delta}$(ILR)w/OT 4.57 0.022